Bob is seven years younger than twice Susan's age. the sum of their ages is 29. How old is each?
Answer by josgarithmetic(39617)

m for bob, f for susan.
m=2f-7 and m+f=29. Various ways to go. (2f-7)+f=29, 2f+f-7=29, 3f-7=29, 3f=29+7, 3f=36, f=12. Do the rest.